The Magellanic Clouds Newsletter

An electronic exchange of Magellanic Clouds information

Edited by: Eva K. Grebel, You-Hua Chu, and Dominik J. Bomans

                                 mcnews@astro.uiuc.edu
                      http://www.astro.uiuc.edu/mcnews/MCNews.html
                         http://www.astro.uni-bonn.de/~mcnews/

No. 18 March 8, 1998

Table of Contents

News  1
Abstracts of 9 refereed papers  2
Abstracts of 3 non-refereed papers 8
Meeting announcement  10

News

Launch of a Dwarf Galaxy Newsletter

On behalf of Elias Brinks (Guanajuato), we are pleased to announce a new electronic newsletter dedicated to dwarf galaxies (excluding the Magellanic Clouds).

The planned newsletter will cover all aspects of dwarf galaxy research, observational as well as theoretical studies, stellar content, ISM, dynamics, formation and evolution, as well as galaxies at low and high redshift. It will contain abstracts of submitted and accepted papers, thesis abstracts, conference announcements, job advertisements, and general news and information.

The dwarf galaxy newsletter has its home at the University of Guanajuato, Mexico at the URL http://www.astro.ugto.mx/~dwarfs/. A European and a US mirror site will also be established. If you wish to subscribe to the dwarf galaxy newsletter please send e-mail to dwarfs@astro.ugto.mx.

Abstracts of Refereed Papers

The C^{18}O/C^{17}O ratio in the Large Magellanic Cloud

A. Heikkilä^{1}, L.E.B. Johansson^{1}, and H. Olofsson^{2}

1 Onsala Space Observatory, S–439 92 Onsala, Sweden 2 Stockholm Observatory, S–133 36 Saltsjöbaden, Sweden

We report detections of J=2–1 line emission from the carbon monoxide isotopomers ^{13}CO, C^{18}O and C^{17}O in the molecular clouds N159W, N113, N44BC, and N214DE in the Large Magellanic Cloud (LMC). ^{13}CO and C^{18}O lines were observed in two additional clouds: N159S in the LMC and N27 in the Small Magellanic Cloud (SMC). While ^{13}CO was detected in both of them, only upper limits to the C^{18}O line emission were obtained. Statistical-equilibrium excitation and radiative transfer calculations were made to infer molecular column densities from the observed line intensities. We estimate an average gas-phase C^{18}O/C^{17}O abundance ratio of 1.6±0.3 in the LMC. This is significantly lower than typical values found in Galactic clouds (by a factor of two) and in centres of starburst galaxies (by a factor of five). We use the C^{18}O/C^{17}O abundance ratio as a measure of the elemental ^{18}O/^{17}O abundance ratio. Provided that current theories of the nucleosynthesis involving ^{17,18}O apply, then the low ^{18}O/^{17}O ratio suggests that massive stars have contributed little to the metal enrichment of the interstellar medium in the LMC in the past. This may be caused by a steep initial mass function (which appears to be the case for field stars in the Magellanic Clouds and in the Galaxy) together with a low average star-formation rate. This explanation contrasts with the present situation in prominent star-formation regions in the LMC, such as 30 Doradus, which form stars at a considerable rate and appear to have initial mass functions similar to star clusters in the Galaxy. The apparent spatial constancy of the ^{18}O/^{17}O abundance ratio, the nominal values for the individual clouds vary between 1.6 and 1.8, indicates a well mixed interstellar medium and/or that the star-formation activity took place globally in the LMC in the past. In the SMC we obtained a lower limit of 17 for the ^{13}CO/C^{18}O ratio (the LMC average is 30), possibly indicating a low ^{18}O abundance here as well. Our data suggests a correlation between the ^{18}O/^{17}O abundance ratio and the metallicity. The high ^{18}O/^{17}O abundance ratio in centres of starburst galaxies could reflect a high metallicity, mainly caused by a high star-formation rate, possibly but not necessarily together with an initial mass function biased towards massive stars.

Accepted by: Astronomy & Astrophysics

For preprints, contact arto@oso.chalmers.se

ASCA X-Ray Spectroscopy of LMC Supernova Remnants and the Metal Abundances of the LMC

J.P. Hughes^{1}, I. Hayashi^{2}, and K. Koyama^{2}

1 Department of Physics and Astronomy, Rutgers University, 136 Frelinghuysen Road, Piscataway, NJ 08854-8019, USA 2 Department of Physics, Graduate School of Science, Kyoto University, Sakyo-ku, Kyoto 606-8502, Japan

We present the results of X-ray spectroscopy of a flux-limited sample of seven middle-aged supernova remnants (SNRs) in the Large Magellanic Cloud (LMC): N23, N49, N63A, DEM71, N132D, 0453−68.5, and N49B. We constructed self-consistent non-equilibrium ionization SNR models assuming a Sedov solution for the dynamical evolution, and then applied the resulting spectral models to the data obtained by the Solid-state Imaging Spectrometer onboard the Advanced Satellite for Cosmology and Astrophysics. All the remnants were reasonably well described by the model, allowing us to derive accurate values for their physical parameters, i.e., ages, densities, initial explosion energies, and metal abundances. The derived explosion energies vary from 5 × 10^{50} erg to 6 × 10^{51} erg. A restricted subset of the sample exists for which the ionization and Sedov dynamical ages agree quite well under the assumption that the electron and ion temperatures are not fully equilibrated at the shock front; for these four SNRs the mean value of the initial explosion energy is (1.1 ± 0.5) × 10^{51} erg. We show that it is likely that the other three remnants exploded within preexisting cavities in the interstellar medium. The limits on high energy X-ray emission (> ∼3 keV) that we present indicate that these SNRs do not contain very luminous pulsar-powered synchrotron nebulae, in general agreement with our picture of them as evolved, middle-aged remnants. We find statistical evidence for enrichment by supernova ejecta in the sense that smaller remnants show a somewhat higher mean metallicity than the larger ones. In the case of DEM71, the putative remnant of a Type Ia supernova, the derived abundance of iron is about a factor of two larger than the other remnants in the sample. These things being said, however, the derived abundances are in general dominated by swept-up interstellar material and so we use the SNR sample to estimate the mean LMC gas-phase abundances. We find that the astrophysically common elements from oxygen to iron are less abundant than the solar values by factors of 2–4. Overall these results are consistent with previous ones based on optical and UV data, but do not show the anomalous overabundance of magnesium and silicon seen by others.

Submitted to: The Astrophysical Journal

For preprints, contact jph@physics.rutgers.edu
Also available from the URL http://www.physics.rutgers.edu/∼jackph/papers

The spin-up of contracting red supergiants

A. Heger^{1} and N. Langer^{2}

1 Max-Planck-Institut für Astrophysik, Karl-Schwarzschild-Str. 1, D–85740 Garching, Germany 2 Institut für Physik, Universität Potsdam, D–14415 Potsdam, Germany

We report on a mechanism which may lead to a spin-up of the surface of a rotating single star leaving the Hayashi line, which is much stronger than the spin-up expected from the mere contraction of the star.

By analyzing rigidly rotating, convective stellar envelopes, we qualitatively work out the mechanism through which these envelopes may be spun up or down by mass loss through their lower or upper boundary, respectively. We find that the first case describes the situation in retreating convective envelopes, which tend to retain most of the angular momentum while becoming less massive, thereby increasing the specific angular momentum in the convection zone and thus in the layers close to the stellar surface. We explore the spin-up mechanism quantitatively in a stellar evolution calculation of a rotating 12 M_{⊙} star, which is found to be spun up to critical rotation after leaving the red supergiant branch.

We discuss implications of this spin-up for the circumstellar matter around several types of stars, i.e., post-AGB stars, B[e] stars, pre-main sequence stars, and, in particular, the progenitor of Supernova 1987A.

Accepted by: Astronomy & Astrophysics

For preprints, contact ahg@mpa-garching.mpg.de
Also available from the URL http://www.mpa-garching.mpg.de/Stars/Preprints/7414.ps.gz

Discovery of a luminous white dwarf in a young star cluster in the Large Magellanic Cloud

Rebecca A.W. Elson, Steinn Sigurdsson, Jarrod Hurley, Melvyn B. Davies & Gerard F. Gilmore

Institute of Astronomy, Madingley Road, Cambridge CB3 0HA, UK

We have identified a candidate ∼ 1 − 2 × 10^{5} year old luminous white dwarf in NGC 1818, a young star cluster in the Large Magellanic Cloud. This discovery strongly constrains the boundary mass M_{c} at which stars stop forming neutron stars and start forming white dwarfs, to M_{c} > ∼ 7.6M_{⊙}.

Submitted to : The Astrophysical Journal, (Letters) For preprints, contact elson@ast.cam.ac.uk

Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9802117

Elemental Abundances for Nova LMC 1990#1

Karen M. Vanlandingham^{1}, Sumner Starrfield^{1}, Steven N. Shore^{2}, and George Sonneborn^{3}

1 Department of Physics and Astronomy, Arizona State University, Tempe, AZ 85287-1504, USA; karenv@suelee.la.asu.edu, sumner.starrfield@asu.edu 2 Department of Physics and Astronomy, Indiana University South Bend, 1700 Mishawaka Ave, South Bend, IN 46634-7111, USA; SShore@paladin.IUSB.edu 3 Laboratory for Astronomy and Solar Physics, Code 681, NASA/GSFC, Greenbelt, MD 20771, USA; sonneborn@stars.gsfc.nasa.gov

We present an analysis of the UV spectra of the fast ONeMg nova LMC 1990#1 Observations with IUE began on 1990 January 18 and continued through 1990 March 15. We find the reddening for the nova to be E(B − V)≃ 0.22. Using the photoionization code CLOUDY 90.03b and the minimization routine MINUIT, we derive elemental abundances for the ejecta. We find the following abundances, relative to solar material: He=1, C=8, N=117, O=20, Ne=62, Mg=10, Al=227, and Si=40. These results are based on independent fits to two different dates of observation, three weeks apart. These abundances are extremely similar to those found for V693 Corona Austrinae 1981. Given the position of V693 CrA and the range of distances found for it, the similar abundances imply that the V693 CrA outburst took place on an older population white dwarf. We find a high N/O ratio which is indicative of an outburst on a massive white dwarf.

Submitted to: Monthly Notices of the Royal Astronomical Society

For preprints, contact karenv@suelee.la.asu.edu
Also available from the URL http://brian.la.asu.edu/~karenv/research.html
or by anonymous ftp at ftp://calvin.physast.uga.edu/pub/preprints/

A radio survey of supersoft, persistent and transient X-ray sources in the Magellanic Clouds

R.P. Fender^{1,2}, K. Southwell^{3}, and A.K. Tzioumis^{4}

1Astronomy Centre, University of Sussex, Falmer, Brighton BN1 9QH, UK 2Astronomical Institute ‘Anton Pannekoek’, University of Amsterdam, and Centre for High Energy Astrophysics, Kruislaan 403, 1098 SJ Amsterdam, The Netherlands 3Department of Astrophysics, Nuclear Physics Building, Keble Road, Oxford OX1 3RH, UK 4Australia Telescope National Facility, CSIRO, PO Box 75, Epping 2121, NSW, Australia

We present a radio survey of X-ray sources in the Large and Small Magellanic clouds with the Australia Telescope Compact Array at 6.3 and 3.5 cm. Specifically, we have observed the fields of five LMC and two SMC supersoft X-ray sources, the X-ray binaries LMC X-1, X-2, X-3 & X-4, the X-ray transient Nova SMC 1992, and the soft gamma-ray repeater SGR 0525-66. None of the targets are detected as point sources at their catalogued positions. In particular, the proposed supersoft jet source RXJ 0513-69 is not detected, placing constraints on its radio luminosity compared to Galactic jet sources. Limits on emission from the black hole candidate systems LMC X-1 and X-3 are consistent with the radio behaviour of persistent Galactic black hole X-ray binaries, and a previous possible radio detection of LMC X-1 is found to almost certainly be due to nearby field sources. The SNR N49 in the field of SGR 0525-66 is mapped at higher resolution than previously, but there is still no evidence for any enhanced emission or disruption of the SNR at the location of the X-ray source.

Accepted by: Monthly Notices of the Royal Astronomical Society For preprints, contact rpf@astro.uva.nl

Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9802180

Optical Gravitational Lensing Experiment.

Distance to the Magellanic Clouds with the Red Clump Stars: Are the Magellanic Clouds 15% Closer than Generally Accepted?

A. Udalski^{1}, M. Szymański^{1}, M. Kubiak^{1}, G. Pietrzyński^{1}, P. Woźniak^{2}, and K. Żebruń^{1}

1 Warsaw University Observatory, Al. Ujazdowskie 4, 00-478 Warszawa, Poland 2 Princeton University Observatory, Princeton, NJ 08544-1001, USA

We present a new distance determination to the Large and Small Magellanic Clouds using the newly developed red clump stars method (Paczyński and Stanek 1998). This new, single-step, Hipparcos calibrated method seems to be one of the most precise techniques of distance determination with very small statistical error due to large number of red clump stars usually available.

The distances were determined independently along four lines-of-sight located at opposite sides of each Magellanic Cloud. The results for each line-of-sight are very consistent. For the SMC we obtain the distance modulus: m − M = 18.56 ± 0.03 ± 0.06 mag (statistical and systematic errors, respectively) and for the LMC: m−M = 18.08±0.03±0.12 mag where systematic errors are mostly due to uncertainty in reddening estimates. Both distances will be refined and systematic errors reduced when accurate reddening maps for our fields are available.

Distance moduli to both Magellanic Clouds are ≈ 0.4 mag smaller than generally accepted values. The modulus to the LMC is in good agreement with the recent determinations from RR Lyrae type stars and upper limit resulting from the SN1987A echo. We suspect that the distance to the LMC and SMC is shorter by about 15% than previously assumed: 42 kpc and 52 kpc, respectively. Calibrations of the period-luminosity relation for Cepheids which give overestimated distances to the LMC and SMC are probably incorrect and require urgent reanalysis.

We also present our color-magnitude diagrams around the red clump for the LMC and SMC. We identify vertical red clump, first noted by Zaritsky and Lin (1997), in the color-magnitude diagram of both Magellanic Clouds and we interpret it as an evolutionary feature rather than unknown stellar population between the LMC and our Galaxy.

Submitted to: Acta Astronomica

For preprints, contact udalski@sirius.astrouw.edu.pl
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9803035

Microlens Parallax Asymmetries Toward the LMC

Andrew Gould

Dept. of Astronomy, Ohio State University, Columbus, OH 43210-1106, USA

If the microlensing events now being detected toward the Large Magellanic Cloud (LMC) are due to lenses in the Milky Way halo, then the events should typically have asymmetries of order 1of the Earth. By contrast, if the lenses are in the LMC, the parallax effects should be negligible. A groundbased search for such parallax asymmetries would therefore clarify the location of the lenses. A modest effort (2 hours per night on a 1 m telescope) could measure 15 parallax asymmetries over 5 years and so marginally discriminate between the halo and the LMC as the source of the lenses. A dedicated 1-m telescope would approximately double the number of measurements and would therefore clearly distinguish between the alternatives. However, compared to satellite parallaxes, the information extracted from ground-based parallaxes is substantially less useful for understanding the nature of the halo lenses (if that is what they are). The backgrounds of asymmetries due to binary-source and binary-lens events are estimated to be approximately 7% and 12% respectively. These complicate the interpretation of detected parallax asymmetries, but not critically.

Submitted to: The Astrophysical Journal

For preprints, contact gould@astronomy.ohio-state.edu
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9802132

Ω_{B} and Ω_{0} From MACHOs and Local Group Dynamics

G. Steigman^{1} and I. Tkachev^{1,2,3}

1Departments of Physics & Astronomy, The Ohio State University, Columbus, OH 43210, USA 2Institute for Nuclear Research of the Academy of Sciences of Russia, Moscow 117312, Russia 3Department of Physics, Purdue University, West Lafayette, IN 47907, USA (Current address)

We obtain restrictions on the universal baryon fraction, f_{B} ≡ Ω_{B}/Ω_{0}, by assuming that the observed microlensing events towards the Large Magellanic Cloud are due to baryonic MACHOs in the halo of the Galaxy and by extracting a bound to the total mass of the Milky Way from the motion of tracer galaxies in the Local Group. We find a lower bound f_{B} > 0.31_{−0.15}^{+0.19}. Consistency with the predictions of primordial nucleosynthesis leads to the further constraint on the total mass density, Ω_{0} < ∼ 0.3_{−0.2}^{+0.1} .

Submitted to: The Astrophysical Journal

Available from the URL http://xxx.lanl.gov/abs/astro-ph/9803008

Abstracts of Non-Refereed Papers

The binary star cluster SL 538/ NGC 2006 (SL 537)

Andrea Dieball^{1} and Eva K. Grebel^{2}

1Sternwarte der Universität Bonn, Auf dem Hügel 71, D-53121 Bonn, Germany 2UCO/Lick Observatory, University of California at Santa Cruz, Santa Cruz, CA 95064, USA

We studied the binary cluster candidate SL 538 / NGC 2006. This cluster pair is located in the northwestern part of the large OB association LH 77 in supergiant shell LMC 4. A third star cluster, KMHK 1019, is located within 5′′ from the cluster pair. We derived ages through isochrone fits and found that the three star clusters have similar ages, which makes near-simultaneous formation likely, possibly all in the same giant molecular cloud. We identified Be star candidates and find the same ratio N(Be)/N(B) for the components of the binary cluster while the amount of Be stars detected in KMHK 1019 and in the surrounding field is considerably lower. Since Be stars are usually rapid rotators this may indicate intrinsically higher rotational velocities in the components of the cluster pair. The IMF derived from the CMDs shows the same slope for both SL 538 and NGC 2006 and is consistent with a Salpeter IMF. We suggest that SL 538 and NGC 2006 are a true binary cluster.

To appear in: Proceedings of the workshop “The Magellanic Clouds and Other Dwarf Galaxies”, Bad Honnef, January 18–22, 1998; eds. T. Richtler & J. Braun.

For preprints, contact adieball@astro.uni-bonn.de

New Distance Determination to the LMC

Nino Panagia ^{1,2}

1 Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA 2 On assignment from the Space Science Department of ESA

Using the new reductions of the IUE light curves by Sonneborn et al. (1997) and an extensive set of HST images of SN 1987A we have repeated and improved Panagia et al.’s (1991) analysis to obtain a better determination of the distance to the supernova.

In this way we have derived an absolute size of the ring R_{abs} = (6.23 ± 0.08) × 10^{17} cm and an angular size R^{”} = 808 ± 17 mas, which give a distance to the supernova d(SN1987A)= 51.4 ± 1.2 kpc and a distance modulus m − M(SN1987A) = 18.55 ± 0.05. Allowing for a displacement of SN 1987A position relative to the LMC center, the distance to the barycenter of the Large Magellanic Cloud is also estimated to be d(LMC) = 52.0 ± 1.3 kpc, which corresponds to a distance modulus of m − M(LMC) = 18.58 ± 0.05.

Revising accordingly the zero point of the Cepheid distance scale, and using the SNIa measured by Sandage and collaborators, one finds a value of the Hubble constant H_{0} = 60 ± 6 km s^{−1} Mpc^{−1}.

Invited Talk at the Workshop “Views on Distance Indicators”, Sant’Agata sui Due Golfi, Italy, 3-6 September, 1997, ed. F. Caputo, Mem. S.A.It., in press. For preprints, contact panagia@stsci.edu

The Recent Star Formation History of the LMC

E.K. Grebel^{1} and W. Brandner^{2,3}

1 UCO/Lick Observatory, University of California at Santa Cruz, Santa Cruz, CA 95064, USA 2 Astronomy Department, University of Illinois, Urbana-Champaign, IL 61801, USA 3 Caltech – JPL/IPAC, Mail Code 100-22, Pasadena, CA 91125, USA

Cepheids and other supergiants are well suited as tracers of the recent star formation history of the LMC. We used 2138 supergiants with UBV photometry from the catalogs of Rousseau et al. (1978), Isserstedt (1979, 1982), Feitzinger & Isserstedt (1983), and Schmidt-Kaler et al. (1998), which we dereddened using the intrinsic UBV color calibration for LMC stars by Isserstedt (1982). We performed an age calibration of UBV colors based on Geneva models by Schaerer et al. (1993).

We also used 1170 Cepheids from Payne-Gaposchkin (1971) and Welch at al. (1995). We calibrated the periods of the Cepheids versus age using Cepheids in 30 open clusters and Geneva isochrones (Grebel & Brandner, in prep.).

The resulting data set covers a time period from approximately 300 million years ago to the present. Our data show the spatial variations in star formation along the LMC bar from the southeast to the northwest and the formation of the currently observed supergiant shells surrounding the bar. We find Isserstedt’s (1984) conclusion confirmed that global star formation occurred stochastically across the face of the LMC. The highly heterogeneous gas distribution prevents self-propagating star formation on a large scale.

The age structure of the LMC does not support the recently suggested scenario for bow-shock-induced star formation (de Boer et al. 1998). Also, the data show that 30 Dor did not form through collision of the supergiant shells LMC 2 and LMC 3. Massive star formation in 30 Dor and LMC 3 started almost simultaneously and spread later toward LMC 2. The supergiant shells that we observe today appear to exhibit several star formation episodes, the oldest dating back 20 to 25 million years. Typical time scales for continuing large-scale star formation on length scales of 0.5 to 1 kpc seem to range from 15 to 30 million years.

To appear in: Proceedings of the workshop “The Magellanic Clouds and Other Dwarf Galaxies”, Bad Honnef, January 18–22, 1998; eds. T. Richtler & J. Braun. For preprints, contact grebel@ucolick.org

Meeting Announcement

Hi in the Local Universe II

Date and venue

The workshop will held at the School of Physics, University of Melbourne, Australia, on September 28–30, 1998. Outline of the scientific program

This is the second in a series of workshops aimed at exploring the scientific issues associated with the distribution of neutral hydrogen in the local universe; from galactic HVCs out to redshifts of 10000 km/s or more. Particular emphasis will be on results from the HIPASS/ZOA survey. Scope of the workshop

The program includes invited papers, contributed papers and poster papers. Limited time is available in the program for oral presentations, however ample provision will be made for poster presentations and for interactive specialist sessions. Requests for oral presentations must be made by the 31st of August 1998. Full audio-visual facilities are available, including video. The proceedings will be published. Scientific Organizing Committee

Mike Disney, Ken Freeman, Trish Henning, Ofer Lahav, Lister Staveley-Smith (co-chair), Rachel Webster (co-chair). Local Organizing Committee

Erwin de Blok, Virginia Kilborn, Mary Putman, Rachel Webster (Chair). Web page / registration

For further information and registration please visit the workshop webpage at

http://astro.ph.unimelb.edu.au/hiconf/, or send e-mail to hiconf@physics.unimelb.edu.au.

Registration for the meeting is now taking place. The form can be downloaded from the workshop webpage. Registration deadline is August 30, 1998.